OpenStudy (anonymous):

f(x)=1-6x solve f(x)=0 f(2x) can be expressed in the form a+bx find the values of a and b

OpenStudy (lalaly):

1-6x=0 1=6x x=1/6

OpenStudy (lalaly):

f(2x)= 1-6(2x) = 1-12x a=1 b=-12